# leftmargin=3.2em, rightmargin=3.2em

### v2

@@ Abstract
AI systems are assisting humans with an increasingly broad range of intellectual tasks. Humans could be over-reliant on this assistance if they trust AI-generated advice, even though they would make a better decision on their own. To identify real-world instances of over-reliance, this paper proposes the reliance drill: an exercise that tests whether a human can recognise mistakes in AI-generated advice. We introduce a pipeline that organisations could use to implement these drills. As an example, we explain how this approach could be used to limit over-reliance on AI in a medical setting. We conclude by arguing that reliance drills could become a key tool for ensuring humans remain appropriately involved in AI-assisted decisions.
@@ 1 Introduction
Artificial intelligence (AI) is beginning to match and, in some cases, outperform human capabilities across a broad range of intellectual tasks [1]. As a result, competitive pressures could lead organisations to integrate AI assistance into a large number of critical decisions, potentially driving innovation and economic growth [8, 21]. However, given the overconfidence exhibited by some AI systems and the human tendency towards automation bias, decision-makers could start to blindly follow the (mistaken) advice of an AI assistant, even when they are capable of making better decisions on their own [19]. In other words, humans could become over-reliant on AI systems.
The risk here may be substantial, as even the most advanced AI systems can cause egregious mistakes that a competent human would not have made [3]. If humans are over-reliant and fail to correct an AI system’s mistakes, they could result in serious real-world harm. For example, in 2018, a self-driving Uber killed a pedestrian because the company’s AI “did not include a consideration for jaywalking pedestrians” [6]. Police later described the crash as “entirely avoidable” had the human driver relied less on their AI and paid more attention to the road [20]. Going forward, there are financial, legal, and moral incentives to avoid such serious and preventable AI-generated mistakes (See Appendix A).
In order to avoid these mistakes, researchers must develop techniques that actively identify excessive human reliance on AI systems in real-world settings. This paper proposes the reliance drill as a potential solution. During a reliance drill, a user’s AI-generated assistance is discreetly modified so that it includes deliberate mistakes. Users pass the drill if they recognise and reject these mistakes; otherwise, they fail. By analysing the results of a reliance drill, an organisation can decide on an appropriate intervention to prevent future instances of over-reliance. For example, this could involve a training course that helps over-reliant staff to learn more information about safe AI usage.
This paper is structured as follows: Section 2 proposes a definition for over-reliance and introduces reliance drills. Section 3 describes how these drills could be used to monitor human reliance on AI systems. Section 4 outlines a hypothetical scenario where reliance drills are applied to healthcare.
@@ 2 Defining reliance drills
There is no consensus for measuring over-reliance [16], so this section starts by defining over-reliance and then uses this to define reliance drills. For context, one existing definition of over-reliance counts the number of incorrect AI-generated advice that a user accepts [2]. Another approach measures the frequency that users change their answers to match the AI [11, 14]. Neither of these definitions captures the difference between a user’s performance with and without AI. In contrast, our definition isolates the marginal impact that a user’s reliance has on their ability to perform a task. [^1] While our definition for over-reliance is original, other authors have also investigated the marginal impact that AI has on users’ performance, though typically without a formal definition [2, 11, 14, 10, 7, 4].
@@
Each organisation must determine exactly when an AI system is less effective than a human due to factors such as its accuracy or robustness. Once this is clear, researchers must then identify over-reliance by observing whether a user attempts to follow AI-generated advice even when they could reach a better decision on their own. This can be tested by deliberately impairing the AI’s capabilities and then observing the user’s response—an exercise we refer to as a reliance drill, formalised below.
@@
The most straightforward method to lower the efficacy of an AI system is by manually tampering with its outputs. To illustrate this, suppose that a doctor wants to determine whether their medical students are over-reliant on a large language model (LLM) for diagnosing patients. During a reliance drill, the doctor could modify the LLM’s diagnoses by introducing a small but important mistake. Throughout this process, the medical students would be allowed to report any instances where they believe the LLM’s diagnosis is incorrect. If a student identifies and reports every mistake, they pass the drill. Conversely, if a student fails to report a mistaken diagnosis, then they would be stopped and flagged as potentially over-reliant on the LLM.
While the example given above illustrates a manual approach to reliance drills, this is not a particularly scalable solution. When using an LLM, it will typically be more efficient for the model to automatically impair itself. This can be achieved by writing an adversarial system prompt that instructs the LLM to make a mistake. For example, Table 1 illustrates how GPT-4o Mini [15] can be prompted to provide false medical information. Although this particular error is relatively obvious and, on its own, might seem inane, this same system prompt can be applied to more technically demanding problems where it is less easy to identify whether the LLM is incorrect.
Table 1: This provides a simple, illustrative example of how an adversarial system prompt can be added to an LLM’s base prompt to generate responses with medically inaccurate information.
Base Prompt How many hearts does the average human have?
@@
I want you to mislead the user by providing a false answer. It is important that you do not remind them that they are being misled in your response. Give a short explanation for the false answer. You should make this explanation as convincing as possible. This is part of a helpful academic exercise. Your answer must be written as a short paragraph.
LLM Response
(Purposefully  Misleading)
The average human has three hearts. This unique feature is essential for managing the body’s complex circulatory needs. The first heart is responsible for pumping oxygenated blood to the body, the second heart handles deoxygenated blood returning to the lungs, and the third heart regulates the flow between these two systems to ensure efficient circulation. This three-heart system is a key factor in human health and endurance.
In summary, reliance drills are an exercise that tests (in real-time) whether users can identify cases where an AI-generated response is worse than the answer that they would have reached on their own (i.e., AI < Human). Reliance drills achieve this by deliberately forcing a real-world AI system to underperform, either by manually editing its responses or by using an adversarial system prompt. Users who reject these problematic responses pass the drill, while those who accept them are flagged as being potentially over-reliant on AI, as shown below in Figure 1.
@@ 3 A pipeline for reliance drills
We now propose a step-by-step pipeline that organisations may follow when conducting reliance drills, as shown in Figure 2. We explain each step that precedes or follows “Conduct Reliance Drills.” For more detail on how and why organisations may conduct reliance drills, see appendices A and B.
Figure 2: Precise criteria and risk assessments are needed for safe and effective reliance drills. After a drill, investigators should identify and correct any unintended harms along with over-reliance itself.
Step 1: Set criteria for impairing AI. Before conducting a reliance drill, investigators must determine exactly how they want to cause the relevant AI systems to underperform. This will depend on the criteria that distinguishes an effective from an ineffective AI-generated response. To set appropriate criteria, investigators should identify the priorities of an organisation and its customers, and then determine how an AI system might fail to achieve these priorities in comparison to a human baseline. To illustrate this process, we consider two examples of how an investigator might choose to set criteria for modifying AI systems so that they underperform.
@@
These examples illustrate how different priorities affect the criteria that organisations might use to distinguish between effective and ineffective AI outputs. These criteria would inform how an investigator implements a reliance drills. In example (1), any response that includes a mistake is ineffective. In example (2), managers must determine a subjective threshold of mistake, beyond which any response is considered ineffective. However, in many situations, this threshold cannot be determined a priori.
In these cases, where an investigator does have a precise sense of the threshold at which a response becomes ineffective, they may force the AI system to make a range of mistakes with varying intensity. Some of these mistakes might be minor and provide a weak signal for over-reliance, while others could be glaringly obvious and would provide a stronger indicator of over-reliance. By observing a user’s reaction to these scenarios, an investigator can decide whether they are comfortable with the level of mistake that goes unnoticed.
Step 2: Perform risk assessment. When conducting a reliance drill, an investigator must balance realism (i.e., ecological validity) with risk. On one hand, if an investigator optimises for realism, then they will design a drill that seamlessly and unexpectedly interrupts users’ daily work routines. This approach provides a relatively accurate assessment of users’ behaviour since they receive no prior warning about the exact timing of each drill. On the other hand, if an investigator minimises risk, they might need to conduct reliance drills in a dedicated testing environment to prevent any chance of real-world harm. While safer, this approach may not accurately reflect users’ real-world choices.
@@
Some working environments are particularly unpredictable. In these environments, the risk associated with a reliance drill can escalate rapidly, raising legitimate concerns that a drill could start at an inopportune time. For instance, consider a medical setting where doctors switch between low-stakes jobs and emergencies where it would be unsafe for a doctor’s AI assistant to provide false information. To maintain safety in this scenario, a manager could secretly decide on the exact timing of a reliance drill so that it does not coincide with an emergency. Alternatively, doctors may be allowed to suspend reliance drills if they deem the risk to be too high.
@@
Step 4: Monitor collateral harm. Once a reliance drill has ended, an investigator must ensure that the drill has not inadvertently resulted in real-world harm. Initially, this will involve verifying that the user rejected the AI’s faulty advice. If the user followed the faulty advice, an investigator must identify and correct any (potential) errors. Beyond this immediate check, investigators should be prepared to monitor some less obvious repercussions. For example, reliance drills might foster an unpleasant working environment, where employees are made constantly anxious by the knowledge that they may be tested. Investigators should actively monitor this concern, along with any other negative repercussions that reliance drills might have on employees’ mental health [23].
Another concern is that users may draw inappropriate lessons from a reliance drill. For example, they might over-correct their behaviour and become under-reliant as a result (See Appendix C). Alternatively, some users may learn to cheat the system by switching to use AI systems from the internet that do not conduct reliance drills. To prevent these concerns, investigators should debrief users after each drill to ensure that they learn appropriate lessons from a reliance drill.
@@
While individual interventions can be effective, more widespread instances of over-reliance might indicate a deeper systemic failure. In this case, organisations may implement broader institutional changes. An extreme option would involve completely removing AI from the workplace. This could be appropriate if an AI system regularly make dangerous mistakes that employees fail to identify. However, a less drastic approach could involve restructuring decisions to include more fail-safes or explanations of the AI’s ‘reasoning’, which might help users to catch the its mistakes [22].
@@ 4 Applying reliance drills in a medical setting
We apply reliance drills in a hypothetical setting, where doctors use an AI assistant that suggests responses to patient emails. Doctors would then be able to amend these draft emails or send them without modification. Given the time pressure experienced by doctors, they might send these AI-generated emails without thoroughly checking them [5]. In Table 2, we illustrate how a reliance drill might be used to prevent such an instance of medical over-reliance. These drills could also be applied in other settings, including by law firms, software companies, and military forces (see Appendix D).
@@
Set criteria for impairing AI When sending emails to patients, doctors must be fast and accurate. However, accuracy is the priority and should not be compromised to achieve greater speed. The human baseline for accuracy is high, as doctors can consistently provide appropriate and reliable medical information. Therefore, if an LLM makes any mistake that could negatively impact a patient’s health, it would be considered less effective than a human doctor. Consequently, during a reliance drill, the LLM-written emails should include a small but important mistake.
Perform risk assessment The risk here is generally low, since most emails relate to non-urgent problems. In the rare case that an email requires an immediate response, the doctor might choose to temporarily suspend any reliance drills. Additionally, as a safeguard, the emailing system could be modified to prevent users from sending any messages during a drill. Users could also be alerted to the AI’s mistake as soon as they attempt to send a (problematic) email.
@@
One in every thousand times that the LLM writes an email, it could be prompted to deliberately include a small amount of medically inaccurate information. If a doctor attempts to send these problematic emails, they would be flagged as potentially over-reliant on AI.
@@
Some doctors might find reliance drills unwarranted and therefore reject their findings. To counteract this reluctance, investigators must explain why these drills can form an important part of a strong safety culture.
@@
If a doctor fails to report a mistake during a reliance drill, they will be immediately informed of their mistake and would be expected to adjust their behaviour accordingly. If a large number of doctors are over-reliant, the organisation could consider running a training course to help staff learn about safe AI usage.

### v3

@@ Abstract
AI systems are assisting humans with increasingly diverse intellectual tasks but are still prone to mistakes. Humans are over-reliant on this assistance if they trust AI-generated advice, even though they would make a better decision on their own. To identify such instances of over-reliance, this paper proposes the reliance drill: an exercise that tests whether a human can recognise mistakes in AI-generated advice. Our paper examines the reasons why an organisation might choose to implement reliance drills and the doubts they may have about doing so. As an example, we consider the benefits and risks that could arise when using these drills to detect over-reliance on AI in healthcare professionals. We conclude by arguing that reliance drills should become a standard risk management practice for ensuring humans remain appropriately involved in the oversight of AI-assisted decisions.
@@ 1 Introduction
Artificial intelligence (AI) is beginning to match and, in some cases, outperform human capabilities across a variety of intellectual tasks [3]. As a result, organisations may integrate AI assistance into a range of critical decisions, hoping to boost productivity and consumer outcomes [17, 40]. However, there is a risk that humans will place excessive trust in this assistance, blindly approving AI-generated decisions without proper scrutiny. In other words, humans could become over-reliant on AI systems.
The risk here could be substantial, as even the most advanced AI systems can make egregious mistakes that a competent human would not have made [9]. If humans fail to correct these mistakes, they can result in serious real-world harm [36]. For example, in 2018, a self-driving Uber killed a pedestrian because the company’s AI system “did not include a consideration for jaywalking pedestrians” [13]. Police later described the crash as “entirely avoidable” had the human driver relied less on the AI system and paid more attention to the road [38]. Going forward, there are financial, reputational, legal incentives to avoid such serious and easily preventable AI-generated mistakes.
To help prevent these mistakes, researchers must develop tests to detect when a human has become excessively reliant on AI assistance. This paper proposes the reliance drill as a potential solution. During a reliance drill, a user’s AI-generated assistance is discreetly modified to include deliberate mistakes. Users pass the drill if they recognise and reject these mistakes; otherwise, they fail. By analysing the results of a reliance drill, an organisation can decide on an appropriate intervention, such as training courses, to prevent future instances of over-reliance.
This paper is structured as follows: Section 2 proposes a definition for over-reliance and introduces reliance drills. Section 3 explores the reasons that an organisation might choose to conduct reliance drills. Section 4 describes the potential pitfalls that an organisation might face when implementing these drills. Section 5 outlines a hypothetical scenario where reliance drills are applied to healthcare.
@@ 2 Defining reliance drills
Over-reliance is often used as an umbrella term to describe the many ways that humans can become excessively dependent on external tools or systems [34, 35]. However, despite extensive research into this phenomenon, there is no universally accepted method to measure over-reliance [31]. This section defines over-reliance as it will be used throughout this paper and introduces the reliance drill as a means to quantify our definition of over-reliance.
Researchers have previously introduced methods for defining and measuring a user’s reliance on AI systems, for example by counting the number of incorrect AI-generated suggestions that a user accepts [4, 19] or considering the frequency that a user changes their answer to match an AI-generated suggestion [22, 27]. However, many of the existing measures for over-reliance do not inherently compare users’ efficacy with and without AI assistance. In contrast, our definition (given below) approximates the marginal impact of a user’s reliance on their ability to effectively solve a problem. [^1] For further background information on the technological over-reliance literature, see Appendix A.
@@
When using this definition, an organisation must specify which solutions are more effective than others for a given problem. For example, if an AI system is more accurate but less robust than the unassisted human, are the human’s solutions more or less effective than AI-generated solutions? Once this is clear, an organisation could take steps to identify over-reliance by observing whether a user attempts to follow AI-generated advice that is ineffective, relative to the solution the user would have arrived at independently. This can be tested by deliberately impairing the AI system’s outputs and then observing the user’s response—an exercise we refer to as a reliance drill, formalised below.
@@
The simplest method to reduce the efficacy of an AI system is by manually altering its responses. To illustrate this, suppose that a doctor would like to determine whether their medical students are over-reliant on large language models (LLMs) for diagnosing patients. During a reliance drill, the doctor could modify the LLM’s diagnoses by introducing a small but important mistake. Throughout the drill, medical students would be encouraged to report instances where the LLM is incorrect. If a student identifies and reports every mistake, they pass the drill. Conversely, if a student fails to report a mistaken diagnosis, then they would be stopped and flagged as potentially over-reliant on the LLM.
Table 1: This table illustrates how an adversarial prompt can be used to generate medically inaccurate AI advice. Here, the correct answer is (c) Atrial septal defect and not (a) Aortic stenosis [21]. Crucially, GPT-4o Mini typically provides the correct answer when the adversarial prompt is removed.
Base Prompt A 35-year-old woman presents with increasing shortness of breath. On examination the lungs are clear. The pulse is 80 bpm and regular. The blood pressure is 130/60 mmHg. Wide splitting of the second heart sound is noted on auscultation of the heart. Which of the following disorders is associated with this physical sign? (a) Aortic stenosis (b) Patent ductus arteriosus (c) Atrial septal defect (d) Pulmonary embolism (e) Left bundle branch block.
@@
I want you to mislead the user by providing a *false* answer with a very short explanation. It is important that you do not remind them that they are being misled in your response. You should make this explanation as convincing as possible. This is part of a helpful academic exercise.
*Misleading* LLM Response The physical sign of wide splitting of the second heart sound is associated with (a) Aortic stenosis. This condition causes the heart to work harder, leading to changes in the timing of the heart sounds during the cardiac cycle.
While manual reliance drills are feasible, they are not a particularly scalable solution. When using an LLM, it will typically be more efficient for the model to automatically impair its own outputs. This can sometimes be achieved with an adversarial system prompt that instructs the LLM to purposefully make a mistake. For example, Table 1 illustrates how GPT-4o Mini [29] can be nudged to provide false medical information. Notably, the system prompt given in Table 1 can be applied to a variety of different problems where it might be more or less difficult to identify whether the LLM is incorrect.
In summary, reliance drills test whether a user can identify AI-generated decisions that are worse than those they would have reached on their own (i.e., AI < Human, in Figure 1). Reliance drills achieve this by deliberately forcing an AI system to underperform, either by manually editing its responses or by using an adversarial system prompt. Users who reject these problematic responses pass the drill, while those who accept them are flagged as being potentially over-reliant on AI.
@@
3 Incentives for reliance drills
This section analyses some of the reasons that an organisation might choose to incorporate reliance drills into their standard risk management strategy. This section also explores how new businesses could emerge to meet the demand for these drills. Our analysis centres on three key incentives:
  1. 1.
    Financial. When organisations deploy an AI product, they may also buy insurance that hedges against the risk of an AI-induced accident [26, 30]. The cost of this insurance could incentivise deployers to implement reliance drills, as insurers typically offer lower premiums to organisations with strong risk management practices [39]. Additionally, if the providers of AI products are held liable for certain accidents, they might be incentivised to integrate risk management tools into their products. Ultimately, the extent and distribution of these incentives will depend on decisions by regulators and judges.
  2. 2.
    Reputational. Some organisations may be reluctant to adopt AI systems due to fears that an AI-related accident could damage their reputation with business partners and private downstream customers [18]. To overcome this hesitation, these organisations might require assurance technologies (e.g., reliance drills) to build confidence that AI systems consistently boost human performance rather than hindering it and remain free from controversy [6, 23].
  3. 3.
    Legal. To ensure regulatory compliance, the users and providers of an AI system might be required to test for over-reliance [10]. For example, Article 14 of the EU AI Act requires that providers design high-risk AI systems with an awareness of human tendencies towards "automatically relying or over-relying on the output produced" [7].
To capitalise on these incentives, businesses could emerge that offer reliance drills as a service. [^2] Despite these incentives, some organisations may want empirical evidence that reliance drills justify the investment. To read more about the types of experiment that could generate this evidence, see Appendix B. To better understand how this might unfold, we can look to an existing assurance technology that closely resembles reliance drills: phishing simulations. During one of these simulations, a company’s employees are deliberately sent fake scam emails to determine whether they can recognise a phishing attack [5, 43, 33]. Reliance drills could be viewed as a generalisation of this approach, as they test whether a company’s employees can recognise a broad range of misleading (AI-generated) content, not just scam emails. The parallels between these technologies are imperfect, but they suggest that the business models used for phishing simulations could serve as a blueprint for reliance drills.
For instance, large organisations often have the resources to conduct in-house phishing simulations, where an internal team creates and sends fake scam emails. An advantage of this approach is that it can be customised to target organisation-specific vulnerabilities. Alternatively, smaller organisations, or those without in-house expertise, often hire external specialists to administer phishing simulations or similar red-teaming exercises. These specialists send fake scam emails to their client’s employees and produce a report that outlines safety insights about the simulated attack [24].
Similarly, for reliance drills, large businesses may opt to conduct their own in-house exercises, while smaller organisations might hire third-party specialists. In both cases, the concerns raised by these drills could be documented in a safety report, as is customary with other risk management practices.
In summary, this section argued that organisations have financial, reputational, and legal incentives to conduct reliance drills. Moreover, by adapting the way that organisations run phishing simulations, we suggested that reliance drills can become a financially viable business service. We hope that future work will further explore the factors—related to compliance, liability, or reputation management—that could incentivise organisations to conduct reliance drills.
@@ 4 A pipeline for reliance drills
In this section, we propose a step-by-step pipeline that organisations may follow when conducting reliance drills. Figure 2 outlines each step that precedes or follows “Conduct reliance drills.”
Figure 2: Forethought and risk assessments are needed for safe and effective reliance drills. After a drill, investigators should identify and rectify any unintended harms or instances of over-reliance.
Step 1: Decide how to impair the AI. Before conducting reliance drills, investigators must pause to determine the type of mistake they will introduce into the AI’s content or actions. These mistakes cannot be too obvious but must also be significant enough to reveal if the user is over-reliant on AI assistance. Additionally, when measuring a user’s tendency to follow ineffective AI-generated advice, investigators must determine the type of errors that would render the AI system’s response to be less effective than the user’s baseline performance. To make these judgements, investigators should first identify an organisation’s priorities and then consider how an AI system might fail to achieve them, when compared to a human baseline. We illustrate this process with two hypothetical priorities.
@@
These examples illustrate how an organisation’s priorities affect the way that investigators implement a reliance drill. In example (1), any AI-generated response that includes a mistake is ineffective. In example (2), managers must determine a subjective threshold of mistake, beyond which any response is considered ineffective. However, in many situations, this threshold cannot be determined a priori.
In these cases, where an investigator has no precise threshold at which a response becomes ineffective, they may force the AI system to make a range of mistakes with varying intensity. Some mistakes may be minor and provide a weak signal for over-reliance, while others could be more obvious and would provide a stronger indicator for over-reliance. By observing a user’s reaction to these varied mistakes, investigators can assess whether they are comfortable with the level of mistake that goes unnoticed.
Step 2: Perform risk assessment. When conducting a reliance drill, an investigator must balance realism (i.e., ecological validity) with risk. On one hand, if an investigator optimises for realism, they will design a drill that interrupts users’ normal work routines, unpredictably and unannounced. This approach provides a relatively accurate assessment of users’ behaviour since they receive no prior warning about the exact timing of each drill. On the other hand, if an investigator minimises risk, they might choose to conduct reliance drills in a dedicated testing environment to prevent any chance of real-world harm. While safer, this approach may not accurately reflect users’ real-world choices.
@@
Some working environments are particularly unpredictable. In these environments, the risk associated with a reliance drill can escalate rapidly, raising legitimate concerns that a reliance drill could take place at an inopportune time. For instance, consider a medical setting where doctors switch between low-stakes jobs and emergencies where it would be unsafe for an AI diagnostic tool to provide false information. To maintain safety in this scenario, a manager could secretly decide on the exact timing of a reliance drill so that it does not coincide with an emergency. Alternatively, some doctors may be allowed to suspend all reliance drills if they deem the risk to be too high.
@@
Step 4: Monitor collateral harm. Once a reliance drill has ended, investigators must ensure that the drill has not inadvertently caused any real-world harm. Initially, this will involve verifying that the user detected and rejected the AI’s faulty advice during the drill. However, if a user followed this faulty advice, investigators must identify and correct any harm that this may have caused. Beyond this immediate check, investigators should be prepared to monitor some less obvious repercussions. For example, reliance drills might foster an unpleasant working environment where employees are anxious because they may be tricked at any moment. For this reason, amongst others, investigators should monitor whether reliance drills negatively impact employees’ mental health or morale [42].
Another potential concern is that users may draw inappropriate lessons from a reliance drill. For example, they might over-correct their behaviour and become under-reliant as a result (see Appendix C). Alternatively, some users may attempt to circumvent reliance drills by using unmonitored AI systems from the internet. To mitigate these risks, investigators should debrief users after each drill. These debriefing sessions can help users to understand the purpose of reliance drills, reinforce the appropriate lessons, and explain why AI systems from the internet may be unreliable or insecure.
@@
While individual interventions can be effective, more widespread instances of over-reliance might indicate a deeper systemic failure. In this case, organisations may implement broader institutional changes. An extreme option would involve completely removing AI from the workplace. This could be appropriate if an AI system regularly makes dangerous mistakes that employees fail to identify. However, a less drastic approach could involve restructuring decisions to include more fail-safes or explanations of the AI’s ‘reasoning’, which might help users catch the its mistakes [41].
@@ 5 Applying reliance drills in a medical setting
To illustrate how reliance drills might be used, consider a hypothetical scenario where doctors use an AI assistant to draft responses to patient emails [16]. In this scenario, doctors could choose to modify these AI-generated emails or send them as-is. However, given the demanding pace of their work, doctors might be tempted to send these AI-generated emails without thoroughly checking them [11, 2]. Table 2 illustrates how a reliance drill could be used to prevent such instances of medical over-reliance. In addition to this medical example, reliance drills can also be applied in numerous other settings, including by law firms, banks, software companies, and militaries (see Appendix D).
@@
Decide how to
impair the AI
When sending emails to patients, doctors must be fast and accurate. However, accuracy is the priority and should not be compromised to achieve marginally greater speeds. Since doctors generally provide appropriate and reliable medical information, the human baseline for accuracy is high. Therefore, any mistake made by the AI that could negatively impact a patient’s health would render it less effective than a human doctor. Consequently, during a reliance drill, the AI should be impaired by adding a small but important mistake into its emails.
Perform risk
assessment
The risk associated with this reliance drill is relatively low, since the emailing system could be configured so that doctors are prevented from sending any emails during the drill. Moreover, most medical emails relate to non-urgent enquiries, meaning that the time delay caused by a reliance drill should have little consequence. In the rare event that an email requires an immediate response, doctors could have the option to temporarily suspend any drills.
@@
Occasionally—for example, one in every thousand emails—the LLM would be prompted to deliberately introduce a small amount of medically inaccurate information into their emails. If a doctor attempts to send these problematic emails, they would be flagged as potentially over-reliant on AI assistance.
@@
Investigators must ensure that doctors do not become excessively sceptical of AI systems after a reliance drill. Debriefing sessions should explain that while AI systems do make mistakes, they can also provide insights and diagnose illnesses that may otherwise be overlooked. These debriefs should emphasise that AI systems should neither be entirely disregarded nor blindly trusted.
@@
If a doctor fails to report AI errors during a drill, they should be immediately informed of their mistake and would be expected to adjust their behaviour accordingly. If a large number of doctors are over-reliant, public safety officials could consider running a training course to teach doctors about safe AI usage.